# EHR-SeqSQL : A Sequential Text-to-SQL Dataset For Interactively Exploring Electronic Health Records

### v1

@@ Abstract
In this paper, we introduce EHR-SeqSQL, a novel sequential text-to-SQL dataset for Electronic Health Record (EHR) databases. EHR-SeqSQL is designed to address critical yet underexplored aspects in text-to-SQL parsing: interactivity, compositionality, and efficiency. To the best of our knowledge, EHR-SeqSQL is not only the largest but also the first medical text-to-SQL dataset benchmark to include sequential and contextual questions. We provide a data split and the new test set designed to assess compositional generalization ability. Our experiments demonstrate the superiority of a multi-turn approach over a single-turn approach in learning compositionality. Additionally, our dataset integrates specially crafted tokens into SQL queries to improve execution efficiency. With EHR-SeqSQL, we aim to bridge the gap between practical needs and academic research in the text-to-SQL domain.
@@ 2.2 Compositional Generalization
Compositional generalization, a major challenge in semantic parsing, enables a model to manage diverse and unfamiliar user queries using components recognized from training. It is typically achieved by creating different training and test splits of interest. Lake and Baroni (2018) discovered that sequence-to-sequence models struggle to learn compositional structures, as demonstrated through their primitive and length splits. More recently, template splits Finegan-Dollak et al. (2018) and TMCD splits Shaw et al. (2020) were proposed to evaluate compositional generalization on text-to-SQL datasets. In this work, we further extend the concept of compositional generalization into an interactive multi-turn setting within the text-to-SQL domain.
@@ 3.2 NLQ Generation
While previous SQL-to-NLQ studies Gan et al. (2022); Wu et al. (2021) decomposed SQL at the clause level for sub-questions and concatenated clauses later, such simple concatenation has a risk of unnaturalness. Thus we create NLQ templates for each corresponding SQL subqueries for the quality of NLQ. To efficiently annotate NLQs for each corresponding SQL subqueries, we first normalize subqueries by replacing specific table names, column names, and condition values to abstract terms such as table and col (see Figure 2). We then manually create NLQ templates for each normalized subqueries. These templates include slots for table names, column names, condition values, SQL functions (e.g. GROUP BY, AVG), and time expressions. Then the actual question is generated from the NLQ templates by the slot-filling process. More details are in Appendix A.3.
@@ 4.2.2 In-context Learning Models
We use ChatGPT for our in-context learning models. Instead of a zero-shot approach, we employ few-shot prompting. We retrieve similar exemplars for each test instance using the BM25 algorithm to use as prompt. For every experiment with in-context learning setting, we use 20-shot. For EHRSQL, which is single-turn, computing the similarity for each question is sufficient. However, for EHR-SeqSQL, which is multi-turn and context-dependent, it’s necessary to consider both the current question and the interaction history. Therefore, we use 10 retrieved examples related to the current question and another 10 related to the entire interaction history. More details such as prompt and few-shot retrieval in our in-context learning baseline are in Appendix E.
@@ 5.1 Generalization to Unseen Interaction Goals
The experimental results are shown in Table 3. In a random split, all models exhibit strong performance with both EHRSQL and EHR-SeqSQL. However, in the compositional split, models typically show superior generalization ability when trained with EHR-SeqSQL. T5 models, in particular, show a significant performance increase, ranging from 7.63%p (from 67.8% to 75.43%), to 30.34%p (from 52.15% to 82.49%). T5-base finetuned in QQ setting has the best generalization ability. It suggests that a task-specific fine-tuning enables effective extraction of necessary information only from NLQs without being distracted by other contextual factors. On the other hand, ChatGPT suffers a performance drop in the QQ setting, which is suspected to be due to the absence of predicted SQL queries within the prompt, which leads ChatGPT to have less information on target representation. Still, a multi-turn setting leads to a significant performance increase in the QS setting. It’s worth noting that ChatGPT has better compositional generalization ability in a single-turn setting than fine-tuned models. Overall, our experiments show that training with EHR-SeqSQL allows the models to generalize well even with unseen questions.
@@ 5.3.1 Effects on Model Performance
As shown in Table 5, the incorporation of special tokens consistently enhances the performance across all settings, regardless of the model variant or the type of interaction history. In the QQ setting, where the absence of prior query history makes the contextual questions more challenging, the special tokens contribute to a substantial performance increase. This is because those tokens simplify referencing previous questions. Specifically, at the question level, the use of these special tokens leads to a performance increase of 14.6%p (from 75.37 to 89.97), and 30%p (from 52.49 to 82.49) at the interaction level. In the QS setting, on the other hand, the reference is more straightforward because the previous SQL information is given. Still, special tokens enhance model performance by reducing the complexity of the target representation. ChatGPT demonstrates a robust performance even without the special tokens in the QS setting. We speculate this is because ChatGPT is trained on diverse data, which likely includes the standard SQL. Therefore, it might be familiar with standard text-to-SQL tasks, which is also consistent with the recent finding Liu et al. (2023).

### v3

@@ Abstract
In this paper, we introduce EHR-SeqSQL, a novel sequential text-to-SQL dataset for Electronic Health Record (EHR) databases. EHR-SeqSQL is designed to address critical yet underexplored aspects in text-to-SQL parsing: interactivity, compositionality, and efficiency. To the best of our knowledge, EHR-SeqSQL is not only the largest but also the first medical text-to-SQL dataset benchmark to include sequential and contextual questions. We provide a data split and the new test set designed to assess compositional generalization ability. Our experiments demonstrate the superiority of a multi-turn approach over a single-turn approach in learning compositionality. Additionally, our dataset integrates specially crafted tokens into SQL queries to improve execution efficiency. With EHR-SeqSQL, we aim to bridge the gap between practical needs and academic research in the text-to-SQL domain. EHR-SeqSQL is available at https://github.com/seonhee99/EHR-SeqSQL.
@@ 2.2 Compositional Generalization
Compositional generalization, a major challenge in semantic parsing, enables a model to manage diverse and unfamiliar user queries using components recognized from training. It is typically achieved by creating different training and test splits of interest. Lake and Baroni (2018) discovered that sequence-to-sequence models struggle to learn compositional structures, as demonstrated through their primitive and length splits. More recently, template splits Finegan-Dollak et al. (2018) and TMCD splits Shaw et al. (2020) were proposed to evaluate compositional generalization on text-to-SQL datasets. Notably, recent efforts Gan et al. (2022a); Liu et al. (2023b) have proposed compositional generalization in multi-turn settings, which aligns with the concepts we use in this paper. However, we aim to extend their approach by proposing and experimenting with two different challenges in multi-turn compositional generalization: unseen (Section 5.1) and longer (Section 5.2) interactions.
@@ 3.2 NLQ Generation
While previous SQL-to-NLQ studies Gan et al. (2022b); Wu et al. (2021) decomposed SQL at the clause level for sub-questions and concatenated clauses later, such simple concatenation has a risk of unnaturalness. Thus we create NLQ templates for each corresponding SQL subqueries for the quality of NLQ. To efficiently annotate NLQs for each corresponding SQL subqueries, we first normalize subqueries by replacing specific table names, column names, and condition values to abstract terms such as table and col (see Figure 2). We then manually create NLQ templates for each normalized subqueries. These templates include slots for table names, column names, condition values, SQL functions (e.g. GROUP BY, AVG), and time expressions. Then the actual question is generated from the NLQ templates by the slot-filling process. More details are in Appendix A.3.
@@ 4.2.2 In-context Learning Models
We use ChatGPT, LLaMA-7B Touvron et al. (2023), and Code-LLaMA-7B Roziere et al. (2023) for our in-context learning models. Instead of a zero-shot approach, we employ few-shot prompting. We retrieve similar exemplars for each test instance using the BM25 algorithm to use as prompt. For every experiment with in-context learning setting, we use 20-shot. For EHRSQL, which is single-turn, computing the similarity for each question is sufficient. However, for EHR-SeqSQL, which is multi-turn and context-dependent, it’s necessary to consider both the current question and the interaction history. Therefore, we use 10 retrieved examples related to the current question and another 10 related to the entire interaction history. More details such as prompt and few-shot retrieval in our in-context learning baseline are in Appendix E.
@@ 5.1 Generalization to Unseen Interaction Goals
The experimental results are shown in Table 3. In a random split, all models exhibit strong performance with both EHRSQL and EHR-SeqSQL. Still, there is a noticeable performance difference between models specifically trained for SQL generation and those that are not, such as LLaMA and Code-LLaMA, which show relatively worse performance.
However, in the compositional split, models generally perform better with the EHR-SeqSQL dataset, which indicates superior generalization ability in multi-turn setting. T5 models, in particular, show a significant performance increase, ranging from 7.63%p (from 67.8% to 75.43%), to 30.34%p (from 52.15% to 82.49%). T5-base finetuned in QQ setting has the best performance. This suggests that a task-specific fine-tuning enables effective extraction of necessary information only from NLQs without being distracted by other contextual factors. Interestingly, T5-3B shows lower performance than T5-base, and this performance gap is particularly pronounced in our proposed compositional split. This might be due to the compositional split having significantly less training data and the evaluation data being perceived as out-of-distribution due to its compositional nature. This result is consistent with previous observations Kale and Rastogi (2020); Zhang et al. (2023); Shaw et al. (2020) which found that smaller models (T5-base or T5-large) can outperform their larger counterpart, T5-3B, outside the general domain or when dealing with out-of-distribution data.
On the other hand, ChatGPT suffers a performance drop in the QQ setting, which is suspected to be due to the absence of predicted SQL queries within the prompt, which leads ChatGPT to have less information on target representation. Still, a multi-turn setting leads to a significant performance increase in the QS setting. It’s worth noting that ChatGPT has better compositional generalization ability in a single-turn setting than fine-tuned models. LLaMA shows very low performance in every setting, indicating it is not well-suited for generating SQL. Code-Llama performs relatively better and demonstrates clearly better performance with our dataset. Overall, our experiments show that training with EHR-SeqSQL allows the models to generalize well even with unseen questions.
@@ 5.3.1 Effects on Model Performance
As shown in Table 5, the incorporation of special tokens consistently enhances the performance across all settings, regardless of the model variant or the type of interaction history. In the QQ setting, where the absence of prior query history makes the contextual questions more challenging, the special tokens contribute to a substantial performance increase. This is because those tokens simplify referencing previous questions. Specifically, at the question level, the use of these special tokens leads to a performance increase of 14.6%p (from 75.37 to 89.97), and 30%p (from 52.49 to 82.49) at the interaction level. In the QS setting, on the other hand, the reference is more straightforward because the previous SQL information is given. Still, special tokens enhance model performance by reducing the complexity of the target representation. ChatGPT demonstrates a robust performance even without the special tokens in the QS setting. We speculate this is because ChatGPT is trained on diverse data, which likely includes the standard SQL. Therefore, it might be familiar with standard text-to-SQL tasks, which is also consistent with the recent finding Liu et al. (2023a).
@@ 6 Conclusion
Acknowledgements
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2019-II190075), National Research Foundation of Korea (NRF) grant (NRF-2020H1D3A2A03100945), and the Korea Health Industry Development Institute (KHIDI) grant (No.HR21C0198), funded by the Korea government (MSIT, MOHW).